# Forget Vectors at Play: Universal Input Perturbations Driving Machine Unlearning in Image Classification

### v2

@@ Abstract
Machine unlearning (MU), which seeks to erase the influence of specific unwanted data from already-trained models, is becoming increasingly vital in model editing, particularly to comply with evolving data regulations like the “right to be forgotten”. Conventional approaches are predominantly model-based, typically requiring retraining or fine-tuning the model’s weights to meet unlearning requirements. In this work, we approach the MU problem from a novel input perturbation-based perspective, where the model weights remain intact throughout the unlearning process. We demonstrate the existence of a proactive input-based unlearning strategy, referred to forget vector, which can be generated as an input-agnostic data perturbation and remains as effective as model-based approximate unlearning approaches. We also explore forget vector arithmetic, whereby multiple class-specific forget vectors are combined through simple operations (e.g., linear combinations) to generate new forget vectors for unseen unlearning tasks, such as forgetting arbitrary subsets across classes. Extensive experiments validate the effectiveness and adaptability of the forget vector, showcasing its competitive performance relative to state-of-the-art model-based methods. Codes are available at https://github.com/Changchangsun/Forget-Vector.
@@ 1 Introduction
To prevent unauthorized use of personal or sensitive data after training and comply with legislation such as the “right to be forgotten” (Hoofnagle et al., 2019), machine unlearning (MU) has garnered increasing attention as a solution to various challenges in vision tasks (Golatkar et al., 2020; Poppi et al., 2023; Warnecke et al., 2021; Fan et al., 2023). In essence, it initiates a reverse learning process to erase the impact of unwanted data (e.g., specific data points, classes, or knowledge concepts) from an already-trained model, while still preserving its utility for information not targeted by an unlearning request. Based on the guarantees provided for data removal from already-trained models, existing MU methods can be broadly categorized into two approaches: exact unlearning (Dong et al., 2024; Guo et al., 2019; Thudi et al., 2022b) and approximate unlearning (Graves et al., 2021; Thudi et al., 2022a; Becker & Liebig, 2022; Izzo et al., 2021). Exact unlearning guarantees the complete and verifiable removal of targeted data, typically achieved by retraining the model from scratch with the data to be forgotten excluded from the training set, a process we refer to as Retrain. However, due to the high computational overhead and lack of scalability of exact unlearning, research has increasingly focused on approximate unlearning methods, which seek to achieve efficient unlearning without requiring full retraining.
@@
Approximate unlearning strikes a balance between computational efficiency and effective data removal, making it practical for many real-world applications. Most existing approximate unlearning techniques are model-based, updating the model’s weights within a limited number of training iterations to eliminate the influence of specific unwanted data, thus avoiding a full retraining process. Representative methods in this category include fine-tuning approaches (Warnecke et al., 2021; Perifanis et al., 2024), gradient ascent techniques (Thudi et al., 2022a; Chen et al., 2024a), and influence function-based methods (Golatkar et al., 2021, 2020). Although these model-based unlearning methods have made significant strides, they often overlook the data-based dimension and its potential impact on MU. For instance, it remains unclear whether current MU approaches generalize effectively to ‘shifted’ forget data, produced by external perturbations. Additionally, the possibility of a data-based MU design that operates without updating model parameters has yet to be explored. Therefore, this work aims to take a step forward in addressing the following research question:
@@
To address (Q), we study MU from a fresh data-based viewpoint: forget vector, a universal input data perturbation designed to promote unlearning effectively; See the schematic overview in Fig. 1. Before developing the forget vector, we explore the rationale for how data-based perturbations complement current model-based MU approaches, as evidenced by these methods’ generalization to common data shifts, including Gaussian noise and adversarial perturbations (Hendrycks & Dietterich, 2019; Goodfellow et al., 2014). Our design for the forget vector is also related to recent input prompting techniques for vision models, known as visual prompting (Bahng et al., 2022b; Chen et al., 2023; Oh et al., 2023) or model reprogramming (Elsayed et al., 2018; Zhang et al., 2022; Chen, 2024), used in transfer learning and model adaptation. These prompting methods learn input perturbations to enable a fixed model to perform well on new tasks, effectively guiding the model to execute tasks it wasn’t originally trained for. From this perspective, our research on the forget vector also explores whether it is possible to append a trainable “prompt” to the input to guide an already-trained neural network in unlearning specific data. We summarize our contributions below.
@@
$\bullet$ We demonstrate the effectiveness of forget vector arithmetic by using class-wise forget vectors to generate new vectors that effectively eliminate the influence of specific data subsets in image classification models.
@@ 2 Related Work
Machine unlearning (MU) in vision has gained significant attention due to the increasing need for privacy preservation, copyright protection, and ethical data removal in machine learning models. Recent studies (Wang et al., 2024b; Pan et al., 2022; Li et al., 2024a; Foster et al., 2024; Lin et al., 2023; Gupta et al., 2021; Kurmanji et al., 2024b; Di et al., 2022; Chen et al., 2024b; Zhang et al., 2024b; Liu et al., 2024a) in this area have primarily focused on two main tasks: image generation and image classification. For image generation, MU techniques (Fan et al., 2023; Li et al., 2024a; Zhang et al., 2024b) have been proposed to prevent models from generating unwanted or harmful content while retaining high-quality outputs. For example, weight saliency methods guide MU by identifying and selectively altering model parameters to eliminate specific content generation (Fan et al., 2023). In image classification, MU methods have explored various approaches to effectively erase certain classes or images from models (Thudi et al., 2022a; Chen et al., 2024a; Golatkar et al., 2020, 2021; Liu et al., 2024a; Pochinkov & Schoots, 2024). Specifically, fine-tuning-based methods update the model incrementally on a modified dataset without the unwanted data points (Warnecke et al., 2021; Perifanis et al., 2024). Gradient ascent-based approaches attempt to reverse the impact of unwanted data by applying gradient ascent to the model parameters (Thudi et al., 2022a; Chen et al., 2024a). Influence-based unlearning techniques estimate the impact of specific data points on model predictions and parameters, and then reverse those contributions to achieve unlearning (Golatkar et al., 2020, 2021). Another line of research investigates the relationship between MU and model pruning, suggesting that model sparsity can help to bridge the gap between approximate and exact unlearning, reducing the need for complex parameter updates (Liu et al., 2024a; Pochinkov & Schoots, 2024). However, most existing MU methods for vision are model-based, often leading to utility degradation after unlearning and incurring high computational costs.
Beyond vision, MU has been applied to other domains, with notable efforts in natural language processing (Shi et al., 2024; Wang et al., 2024a, 2023a; Liu et al., 2024b), graph-based data (Li et al., 2024b; Dong et al., 2024), and time-series data (Du et al., 2019).
@@
This approach has gained attention as an effective method to modify or repurpose pre-trained models for new tasks or specific objectives without the need for full retraining. It is particularly beneficial for reducing computational costs and leveraging the existing knowledge within models. Key techniques in model adaptation include: Visual prompting (Jia et al., 2022; Wang et al., 2023b; Liu et al., 2023; Hossain et al., 2024; Zhang et al., 2024a) maintains the pre-trained model’s parameters fixed and adapts the input to enable the model to perform different tasks. For example, introducing trainable parameters in the input space while keeping the model backbone frozen can achieve comparable results with reduced computational overhead. Model reprogramming (Tsai et al., 2020; Elsayed et al., 2018; Yang et al., 2021; Melnyk et al., 2023) involves keeping a pre-trained model unchanged while modifying its inputs to adapt the model for new tasks. For example, adversarial perturbations can be applied to inputs at test time, allowing the model to perform a specific task dictated by the perturbations, even if that task was not originally intended for the model. Feature-based domain adaptation (Wang et al., 2018; Tahmoresnezhad & Hashemi, 2017) applies transformations or mapping techniques to the input data, aligning the feature distributions between the source and target domains while keeping the model unchanged.
@@ 3 Preliminaries on MU and Problem Statement
In this section, we introduce the fundamentals of machine unlearning (MU), including its formulation, commonly-used methods, and evaluation metrics, and motivate our focus: an data-based forget vector design for achieving MU.
@@
In this work, we focus on the problem of MU for image classification. Let $\mathcal{D}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N}$ represent a training set with $N$ examples, where $\mathbf{x}_{i}$ denotes the $i$th image data, and $y_{i}$ denotes its corresponding class label. We introduce a forget set $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$, which specifies the training samples targeted for unlearning. Accordingly, the complement of $\mathcal{D_{\mathrm{f}}}$ is the retain set, i.e., $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D_{\mathrm{f}}}$. The goal of MU is to efficiently and effectively eliminate the influence of $\mathcal{D_{\mathrm{f}}}$ on an already-trained model ${{\bm{\theta}}}$, so that the performance of the post-unlearning model closely approximates that of a model retrained from scratch on $\mathcal{D_{\mathrm{r}}}$ (i.e., excluding the impact of $\mathcal{D_{\mathrm{f}}}$ from scratch). Therefore, such a retraining method (referred to as Retrain) is typically considered as the glad standard of MU (Thudi et al., 2022a; Jia et al., 2023). However, since Retrain is computationally intensive, most popular MU approaches instead address an unlearning optimization problem using the forget and retain sets to update the model parameters ${{\bm{\theta}}}$, starting from the originally pre-trained model, denoted by ${{\bm{\theta}}_{\mathrm{o}}}$. This yields the following optimization problem for MU:
@@
Model Update-based MU Methods and Evaluation.
@@
In what follows, we introduce several representative MU approaches that serve as approximations to Retrain. (a) Fine-tuning (FT) (Warnecke et al., 2021): This approach treats the MU problem as a continual learning task, defining the unlearning objective $\ell_{\mathrm{MU}}$ as a training objective that fine-tunes ${\bm{\theta}}$ over $\mathcal{D}_{\mathrm{r}}$ to induce catastrophic forgetting of $\mathcal{D}_{\mathrm{f}}$. (b) Random labeling (RL) (Golatkar et al., 2020): This approach specifies the unlearning objective $\ell_{\mathrm{MU}}$ by assigning random labels or features to the data in $\mathcal{D}_{\mathrm{f}}$, thereby enforcing model forgetting. (c) Gradient ascent (GA) (Thudi et al., 2022a): This approach employs the negative of the FT loss to reverse the training impact associated with the data in $\mathcal{D}_{\mathrm{f}}$. (d) Localization-informed unlearning (Jia et al., 2023; Fan et al., 2023): This method identifies a subset of model weights critical to the unlearning task (e.g., through model sparsity (Jia et al., 2023) or gradient saliency (Fan et al., 2023)) and incorporates this weight localization as a prior to solve the unlearning problem in (2).
Given an unlearned model (denoted as ${{\bm{\theta}}_{\mathrm{u}}}$) after solving (2), unlearning performance is evaluated in two main areas: unlearning effectiveness, which measures whether the target data/information has been successfully removed, and utility retention, which assesses whether unlearning has preserved the model’s classification ability on unaffected data. Following the evaluation benchmark in (Jia et al., 2023), unlearning effectiveness can be quantified by two metrics: unlearning accuracy (UA), defined as $1-$the model’s accuracy on $\mathcal{D_{\mathrm{f}}}$ (higher UA indicates better unlearning), and membership inference attack performance on $\mathcal{D_{\mathrm{f}}}$, termed MIA-Efficacy, where higher prediction accuracy on non-training samples indicates better unlearning. Utility retention is measured by retain accuracy (RA), reflecting the model’s accuracy on $\mathcal{D_{\mathrm{r}}}$, and testing accuracy (TA), which is the accuracy on the original test set.
@@
where $\bm{\delta}$ is the perturbation variable, applied linearly to the forget samples as $\mathbf{x}^{\prime}\mathrel{\mathop{:}}=\mathbf{x}+\bm{\delta}$ for $\mathbf{x}\in\mathcal{D_{\mathrm{f}}}$, similar to visual prompting (Bahng et al., 2022a) and adversarial examples (Goodfellow et al., 2014). Unlike model-based MU methods, we will detail the unlearning objective function required for designing the forget vector in our later method sections.
To address (4), we are motivated to explore two research questions: (Q1) How do “perturbations” applied to forget data affect unlearning performance? (Q2) How can we effectively design the forget vector $\bm{\delta}$ to solve problem (4)? These two questions are interconnected: the answer to (Q1) offers a sensitivity analysis of MU to data shifts within the forget set, guiding how the specific shift induced by the forget vector can be optimized for effective unlearning in (Q2). Therefore, the following Secs. 4-5 address (Q1) and (Q2) in sequence. For (Q1), the next section analyzes performance through an evaluation lens on a given unlearned model, using data perturbations applied via standard data augmentation techniques or adversarial perturbations.

### v4

@@ Abstract
Machine unlearning (MU), which seeks to erase the influence of specific unwanted data from already-trained models, is becoming increasingly vital in model editing, particularly to comply with evolving data regulations like the “right to be forgotten”. Conventional approaches are predominantly model-based, typically requiring retraining or fine-tuning the model’s weights to meet unlearning requirements. In this work, we approach the MU problem from an input perturbation-based perspective, where the model weights remain intact throughout the unlearning process. We demonstrate the existence of a proactive input-based unlearning strategy, referred to forget vector, which can be generated as an input-agnostic data perturbation and remains as effective as model-based approximate unlearning approaches. We also explore forget vector arithmetic, whereby multiple class-specific forget vectors can be combined through simple operations (e.g., linear combinations) to generate new forget vectors for unseen unlearning tasks, such as forgetting arbitrary subsets across classes. Extensive experiments validate the effectiveness and adaptability of the forget vector, showcasing its competitive performance relative to state-of-the-art model-based methods while achieving superior parameter efficiency. Codes are available at https://github.com/Changchangsun/Forget-Vector.
@@ 1 Introduction
To prevent unauthorized use of personal or sensitive data after training and comply with legislation such as the “right to be forgotten” (Hoofnagle et al., 2019), machine unlearning (MU) has garnered increasing attention as a solution to various challenges in vision tasks (Golatkar et al., 2020; Poppi et al., 2023; Warnecke et al., 2021; Fan et al., 2023). In essence, it initiates a reverse learning process to erase the impact of unwanted data (e.g., specific data points, classes, or knowledge) from an already-trained model, while still preserving its utility for information not targeted by an unlearning request. Based on the guarantees provided for data removal from already-trained models, existing MU methods can be broadly categorized into two approaches: exact unlearning (Dong et al., 2024; Guo et al., 2019; Thudi et al., 2022b) and approximate unlearning (Graves et al., 2021; Thudi et al., 2022a; Becker & Liebig, 2022; Izzo et al., 2021). The former guarantees the complete and verifiable removal of targeted data, typically achieved by retraining the model from scratch with the data to be forgotten excluded from the training set, a process we refer to as Retrain. However, due to the high computational overhead, research has increasingly focused on approximate unlearning methods, which seek to achieve efficient unlearning without requiring full retraining.
@@
Approximate unlearning strikes a balance between computational efficiency and effective data removal, making it practical for many real-world applications. Most existing approximate unlearning techniques are model-based, updating the model’s weights within a limited number of training iterations to eliminate the influence of specific unwanted data, thus avoiding a full retraining process. Representative methods in this category include fine-tuning approaches (Warnecke et al., 2021; Perifanis et al., 2024), gradient ascent techniques (Thudi et al., 2022a; Chen et al., 2024a), and influence function-based methods (Golatkar et al., 2021, 2020).
Although the model-based unlearning methods have made significant strides, they often overlook the data-based dimension and its potential impact on MU. For instance, it remains unclear whether current MU approaches generalize effectively to ‘shifted’ forget data. Additionally, the possibility of a data-based MU design that operates without updating model parameters has yet to be explored. Therefore, we ask:
@@
To address (Q), we study MU from a fresh data-based viewpoint: forget vector, a universal input data perturbation designed to promote unlearning effectively; See the schematic overview in Fig. 1. Before developing the forget vector, we explore the rationale for how data perturbations complement current model-based MU approaches, as evidenced by these methods’ generalization to common data shifts, including Gaussian noise and adversarial perturbations (Hendrycks & Dietterich, 2019; Goodfellow et al., 2014). To design the forget vector, we draw inspiration from recent input prompting techniques for vision models, known as visual prompting (Bahng et al., 2022b; Chen et al., 2023; Oh et al., 2023) or model reprogramming (Elsayed et al., 2018; Zhang et al., 2022; Chen, 2024), used in transfer learning and model adaptation. These prompting methods learn input perturbations to enable a fixed model to perform well on new tasks, effectively guiding the model to execute tasks it wasn’t originally trained for. From this perspective, our research on the forget vector also explores whether it is possible to append a trainable “prompt” to the input to guide an already-trained neural network in unlearning specific data. The proposed forget vector allows the unlearner to modify user inputs targeted for deletion, offering a flexible and efficient approach to unlearning while potentially achieving significant parameter efficiency. We summarize our contributions below.
@@
$\bullet$ We demonstrate the effectiveness of forget vector arithmetic by using precomputed class-wise forget vectors to generate new vectors that effectively eliminate the influence of specific data subsets in image classification models, e.g., in the scenario of random data forgetting.
@@ 2 Related Work
Machine unlearning (MU) in vision has gained significant attention due to the increasing need for privacy preservation, copyright protection, and ethical data removal in machine learning models. Recent studies (Wang et al., 2024b; Pan et al., 2022; Li et al., 2024a; Foster et al., 2024b; Lin et al., 2023; Gupta et al., 2021; Kurmanji et al., 2024b; Di et al., 2022; Chen et al., 2024b; Zhang et al., 2024c; Liu et al., 2024a) in this area have primarily focused on two main applications: image classification and image generation.
In image classification, MU methods have explored various ways to erase certain classes or images from models (Thudi et al., 2022a; Chen et al., 2024a; Golatkar et al., 2020, 2021; Liu et al., 2024a; Pochinkov & Schoots, 2024). Specifically, fine-tuning-based methods update the model incrementally on a modified dataset without the unwanted data points (Warnecke et al., 2021; Perifanis et al., 2024). Gradient ascent-based approaches attempt to reverse the impact of unwanted data by applying gradient ascent to model parameters (Thudi et al., 2022a; Chen et al., 2024a). Influence-based unlearning techniques estimate and negate the effect of specific data points on model predictions and parameters to achieve unlearning (Golatkar et al., 2020, 2021). Another line of research explores the relationship between MU and model pruning, suggesting that model sparsity can help to bridge the gap between approximate and exact unlearning, reducing the need for complex parameter updates (Liu et al., 2024a; Pochinkov & Schoots, 2024).
For image generation, MU techniques (Fan et al., 2023; Li et al., 2024a; Zhang et al., 2024c) have been proposed to prevent models from generating unwanted or harmful content while retaining high-quality outputs. For example, weight saliency methods guide MU by identifying and selectively altering model parameters to eliminate specific content generation (Fan et al., 2023). Beyond vision, MU has been applied to other domains, with notable efforts in natural language processing (Shi et al., 2024; Wang et al., 2024a, 2023a; Liu et al., 2024b), graph-based data (Li et al., 2024b; Dong et al., 2024), and time-series data (Du et al., 2019). However, most existing MU methods are model-based, requiring updates to model parameters and consequently incurring high computational costs.
@@
This approach aims to modify or repurpose pre-trained models for new tasks or specific objectives without the need for full retraining. It is particularly beneficial for reducing computational costs and leveraging existing knowledge within models. Key techniques in model adaptation include: Visual prompting (Jia et al., 2022; Wang et al., 2023b; Liu et al., 2023; Hossain et al., 2024; Zhang et al., 2024a) maintains the pre-trained model’s parameters fixed and adapts the input to enable the model to perform different tasks. For example, introducing trainable parameters in the input space while keeping the model backbone frozen can achieve comparable results with reduced computational overhead. Model reprogramming (Tsai et al., 2020; Elsayed et al., 2018; Yang et al., 2021; Melnyk et al., 2023) involves keeping a pre-trained model unchanged while modifying its inputs to adapt the model for new tasks. For example, adversarial perturbations can be applied to inputs at test time, allowing the model to perform a specific task dictated by the perturbations, even if that task was not originally intended for the model. Feature-based domain adaptation (Wang et al., 2018; Tahmoresnezhad & Hashemi, 2017) applies transformations or mapping techniques to the input data, aligning the feature distributions between the source and target domains while keeping the model unchanged.
@@ 3 Preliminaries on MU and Problem Statement
In this section, we introduce the fundamentals of machine unlearning (MU), including its formulation, commonly-used methods, evaluation metrics, and motivate our focus: a data-based forget vector design for achieving MU.
@@
In this work, we focus on the problem of MU for image classification. Let $\mathcal{D}=\{\mathbf{x}_{i},y_{i}\}_{i=1}^{N}$ represent a training set with $N$ examples, where $\mathbf{x}_{i}$ denotes the $i$th image data, and $y_{i}$ denotes its corresponding class label. Following the classic MU setting (Golatkar et al., 2020; Kurmanji et al., 2024a, b; Fan et al., 2023), we introduce a forget set $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$, which specifies the training samples targeted for unlearning. Accordingly, the complement of $\mathcal{D_{\mathrm{f}}}$ is the retain set, i.e., $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D_{\mathrm{f}}}$. The goal of MU is to efficiently and effectively eliminate the influence of $\mathcal{D_{\mathrm{f}}}$ on an already-trained model ${{\bm{\theta}}_{\mathrm{o}}}$, so that the performance of the post-unlearning model closely approximates that of a model retrained from scratch on $\mathcal{D_{\mathrm{r}}}$ (i.e., excluding the impact of $\mathcal{D_{\mathrm{f}}}$ from scratch). Therefore, such a retraining method (referred to as Retrain) is typically considered as the gold standard of MU (Thudi et al., 2022a; Jia et al., 2023). However, since Retrain is computationally intensive, most popular MU approaches instead address an unlearning optimization problem using the forget and retain sets to update the model parameters ${{\bm{\theta}}}$, starting from the originally pre-trained model ${{\bm{\theta}}_{\mathrm{o}}}$. This yields the following optimization problem for MU:
@@
Model-based MU Methods and Evaluation.
@@
In what follows, we introduce several representative MU approaches that serve as approximations to Retrain. (a) Fine-tuning (FT) (Warnecke et al., 2021): This approach treats the MU problem as a continual learning task, defining the unlearning objective $\ell_{\mathrm{MU}}$ as a training objective that fine-tunes ${{\bm{\theta}}_{\mathrm{o}}}$ over $\mathcal{D}_{\mathrm{r}}$ to induce catastrophic forgetting of $\mathcal{D}_{\mathrm{f}}$. (b) Random labeling (RL) (Golatkar et al., 2020): This approach specifies the unlearning objective $\ell_{\mathrm{MU}}$ by assigning random labels or features to the data in $\mathcal{D}_{\mathrm{f}}$, thereby enforcing model forgetting. (c) Gradient ascent (GA) (Thudi et al., 2022a): This approach employs the negative of the FT loss to reverse the training impact associated with the data in $\mathcal{D}_{\mathrm{f}}$. (d) Localization-informed unlearning (Jia et al., 2023; Fan et al., 2023): This method identifies a subset of model weights critical to the unlearning task (e.g., through model sparsity (Jia et al., 2023) or gradient saliency (Fan et al., 2023)) and incorporates this weight localization as a prior to solve the unlearning problem in (2).
Given an unlearned model (denoted as ${{\bm{\theta}}_{\mathrm{u}}}$) after solving (2), unlearning performance is evaluated in two main areas: unlearning effectiveness, which measures whether the target data/information has been successfully removed, and utility retention, which assesses whether unlearning has preserved the model’s classification ability on unaffected data. Following the evaluation pipeline in (Jia et al., 2023), unlearning effectiveness is quantified by two metrics: unlearning accuracy (UA), defined as $1-$the model’s accuracy on $\mathcal{D_{\mathrm{f}}}$ (higher UA indicates better unlearning), and membership inference attack performance on $\mathcal{D_{\mathrm{f}}}$, termed MIA-Efficacy, where higher prediction accuracy on non-training samples indicates better unlearning (see Appendix A). Utility retention is measured by retain accuracy (RA), reflecting the model’s accuracy on $\mathcal{D_{\mathrm{r}}}$, and testing accuracy (TA), which is the accuracy on the original test set. Notably, TA is assessed on the entire original test set, except in the case of class-wise forgetting, where test samples from the forgotten class are excluded from evaluation.
@@
where $\bm{\delta}$ is the perturbation variable, applied linearly to the forget samples as $\mathbf{x}^{\prime}\mathrel{\mathop{:}}=\mathbf{x}+\bm{\delta}$ for $\mathbf{x}\in\mathcal{D_{\mathrm{f}}}$, similar to visual prompting (Bahng et al., 2022a) and adversarial examples (Goodfellow et al., 2014). In practice, since the model remains unchanged, the unlearner can compute the forget vector based on the forget request (forget set) and append it to model inputs to process user-initiated unlearning requests. In this work, we do not consider counter-unlearning adversaries that intentionally negate the effect of the forget vector. We will detail the unlearning objective function required for designing the forget vector in our later method sections.
Based on (4), we are motivated to explore two research questions: (Q1) How do “perturbations” applied to forget data affect unlearning performance? (Q2) How can we effectively design the forget vector $\bm{\delta}$ to solve problem (4)? These two questions are interconnected: the answer to (Q1) offers a sensitivity analysis of MU to data shifts within the forget set, guiding how the specific shift induced by the forget vector can be optimized for effective unlearning in (Q2). Therefore, the following Secs. 4-5 address (Q1) and (Q2) in sequence. For (Q1), the next section analyzes performance through an evaluation lens on a given unlearned model, using data perturbations applied via standard data augmentation operations or adversarial perturbations.